# DefVINS: Visual–Inertial Odometry for Deformable Scenes

### v1

@@ Abstract
Deformable scenes violate the rigidity assumptions underpinning classical visual–inertial odometry (VIO), often leading to over–fitting to local non–rigid motion or severe drift when deformation dominates visual parallax. We introduce DefVINS, a visual–inertial odometry framework that explicitly separates a rigid, IMU–anchored state from a non–rigid warp represented by an embedded deformation graph. The system is initialized using a standard VIO procedure that fixes gravity, velocity, and IMU biases, after which non–rigid degrees of freedom are activated progressively as the estimation becomes well conditioned. An observability analysis is included to characterize how inertial measurements constrain the rigid motion and render otherwise unobservable modes identifiable in the presence of deformation. This analysis motivates the use of IMU anchoring and informs a conditioning–based activation strategy that prevents ill–posed updates under poor excitation. Ablation studies demonstrate the benefits of combining inertial constraints with observability–aware deformation activation, resulting in improved robustness under non–rigid environments. Source will be released upon acceptance.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM), and its core state estimation component, Visual–Inertial Odometry (VIO), are foundational technologies driving modern mobile robotics and Augmented Reality (AR). Visual odometry and SLAM is a very active field of research with an abundance of applications in these fields, demonstrating remarkable maturity through various sensor configurations, including monocular [1] and stereo [2].
@@
Among these, the fusion of vision with an Inertial Measurement Unit (IMU) stands out as a particularly robust solution. Adding an IMU helps dealing with untextured environments and rapid motions and makes roll and pitch directly observable [3]. The IMU, which provides high–frequency measurements of linear acceleration and angular velocity, effectively anchors the system’s short–term pose dynamics and provides crucial observability to the orientation axes unobservable in pure monocular vision. On the other hand, the camera complements the IMU with external referencing to the environment in 6 Degrees of Freedom (DoF), correcting the inevitable drift resulting from IMU sensor biases over time [4, 5]. State–of–the–art VIO systems, such as VINS–Mono [5] and OKVIS [6], have proven highly effective in static, rigid environments, establishing them as the de facto standard for ego–motion estimation.
@@
However, the efficacy of classical VIO hinges entirely upon the fundamental assumption of scene rigidity [7]. This assumption is severely violated in scenarios featuring deformable objects such as human bodies or clothing. When the geometric model fails to account for non–rigid motion, VIO systems experience significant issues: the estimation often suffers from early over–fitting of the rigid model to local non–rigid motion, or worse, considerable drift when deformation parallax dominates inter–frame motion [8]. Consequently, achieving robust and accurate localization in truly dynamic and deformable scenes remains an open and critical research problem [9].
@@
Beyond modeling challenges, the transition from rigid to deformable environments fundamentally alters the observability properties of the estimation problem. While classical analyses of VIO observability have shown that inertial sensing renders scale, gravity direction, and roll–pitch observable in rigid scenes, the impact of these properties in the presence of non–rigid motion remains largely underexplored. Deformation introduces additional latent degrees of freedom that can strongly couple with camera motion, leading to severe ill–conditioning and ambiguity when relying on visual information alone. In this context, inertial measurements play a critical role by anchoring the rigid body dynamics over short time horizons, effectively constraining the solution space and mitigating spurious correlations between ego–motion and non–rigid deformation. This work therefore includes an explicit observability analysis of the visual–inertial deformable odometry problem, highlighting how inertial terms significantly improve conditioning and provide essential structural constraints that are otherwise absent in purely visual formulations.
To address the instability and ill-conditioning inherent in extending VIO to non-rigid scenes, we introduce DefVINS: Robust Visual–Inertial Odometry for Deformable Scenes. Our framework utilizes an embedded deformation graph to explicitly model the non-rigid warp, carefully separating the rigid, IMU-anchored state from the scene’s non-rigid degrees of freedom.
@@ II Related Work
Visual–Inertial Odometry (VIO) is a well–established paradigm for robust ego–motion estimation in rigid environments through the tight fusion of visual and inertial measurements [10, 11, 12]. Inertial sensing provides high–rate motion constraints that resolve metric scale, stabilize roll and pitch, and mitigate drift, while visual observations offer global referencing. The observability and consistency properties of such systems have been studied extensively under rigid–scene assumptions, showing that scale, gravity, velocity, and IMU biases become observable only under sufficient excitation and appropriate modeling choices [13, 14, 15]. These analyses, however, fundamentally assume that all observed features belong to a single rigid body.
When this assumption is violated, visual residuals become inconsistent and may bias the estimation or lead to divergence. Several approaches address dynamic scenes by detecting and suppressing independently moving elements through semantic segmentation or motion consistency checks [16]. While effective under moderate dynamics, these methods do not explicitly model deformation and instead discard measurements, limiting their applicability in scenes dominated by non–rigid motion. Other dynamic SLAM pipeline, such as Detect–SLAM, focus on handling multiple moving objects but remain purely visual and do not address continuous deformation or metric consistency, as summarized in Table I.
A complementary line of work explicitly models non–rigid scene geometry. DynamicFusion [17], DefSLAM [18], and NR–SLAM [19] represent deformation using embedded deformation graphs or learned warp fields, enabling compelling reconstructions. However, these systems rely exclusively on visual cues, lack inertial anchoring, and do not preserve metric scale, resulting in a strong coupling between camera motion and scene deformation and, consequently, ill–conditioned pose estimation under strong non–rigid motion.
Method Def. IMU Scale Obs. Analysis
ROVIO [20]
SVO2 [21]
VINS–Mono [11]
OKVIS [10]
ORB–SLAM3 [12]
DynaSLAM [16] (✗)
Detect–SLAM [22] (✗)
CUDA–SIFT–SLAM [23]
DynamicFusion [17]
DefSLAM [18]
NR–SLAM [19]
DefVINS (Ours)
TABLE I: Comparison with representative SLAM and VIO approaches. The table highlights whether each method explicitly models non–rigid deformation (Def.), uses inertial sensing for metric state estimation (IMU / Scale), and provides an explicit observability or conditioning analysis of the estimation problem (Obs. Analysis).
@@
Table I summarizes representative SLAM and visual–inertial approaches. To the best of our knowledge, no existing method jointly (i) integrates an explicit deformation model within a visual–inertial odometry pipeline, (ii) maintains a rigid, IMU–anchored reference to ensure metric consistency, and (iii) provides an explicit observability or conditioning analysis of the resulting non–rigid estimation problem. We emphasize that the Obs. Analysis column refers to an explicit characterization of observability or conditioning of the proposed estimation model, rather than the implicit observability properties of a given sensor configuration.
@@
DefVINS addresses these limitations by embedding a deformation graph within a visual–inertial optimization framework and by explicitly analyzing the observability properties of the resulting system, demonstrating that inertial constraints significantly improve conditioning even under strong non–rigid motion.
III Notation and Preliminaries
Vectors are denoted by bold lowercase letters ($\mathbf{x}$), unit vectors by a check accent ($\check{\mathbf{x}}$), and matrices by bold uppercase letters ($\mathbf{A}$). Scalars are represented by lowercase letters ($a$). Rotation matrices $\mathtt{R}_{ab}\in\mathrm{SO}(3)$ denote the orientation of frame $b$ with respect to frame $a$; when the reference frame is the world frame $w$, the first subscript is omitted, i.e., $\mathtt{R}_{a}\doteq\mathtt{R}_{wa}$.
IMU kinematic propagation between time instants $t_{i}$ and $t_{j}$ follows the standard on-manifold preintegration formulation [26] and is given by
@@ III Notation and Preliminaries
$\mathtt{R}_{j}$ $=\mathtt{R}_{i}\prod_{k=i}^{j-1}\mathrm{Exp}\!\left(\big(\tilde{\boldsymbol{\omega}}_{k}-\mathbf{b}^{g}_{k}-\boldsymbol{\eta}^{gd}_{k}\big)\,\Delta t\right),$ (1)
$\mathbf{v}_{j}$ $=\mathbf{v}_{i}+\mathbf{g}\,\Delta t_{ij}+\sum_{k=i}^{j-1}\mathtt{R}_{k}\big(\tilde{\mathbf{a}}_{k}-\mathbf{b}^{a}_{k}-\boldsymbol{\eta}^{ad}_{k}\big)\,\Delta t,$ (2)
$\mathbf{t}_{j}$ $=\mathbf{t}_{i}+\sum_{k=i}^{j-1}\Big[\mathbf{v}_{k}\Delta t+\tfrac{1}{2}\mathbf{g}\,\Delta t^{2}+\tfrac{1}{2}\mathtt{R}_{k}\big(\tilde{\mathbf{a}}_{k}-\mathbf{b}^{a}_{k}-\boldsymbol{\eta}^{ad}_{k}\big)\,\Delta t^{2}\Big],$ (3)
where $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ denote the linear velocities at times $t_{i}$ and $t_{j}$, $\mathbf{t}_{i}$ and $\mathbf{t}_{j}$ the corresponding positions, $\mathbf{g}$ the gravity vector, and $\Delta t$ the IMU sampling period, with $\Delta t_{ij}$ denoting the total integration interval. The measurements $\tilde{\boldsymbol{\omega}}_{k}$ and $\tilde{\mathbf{a}}_{k}$ correspond to the gyroscope and accelerometer readings at time $k$, affected by biases $\mathbf{b}^{g}_{k}$, $\mathbf{b}^{a}_{k}$ and additive noise terms $\boldsymbol{\eta}^{gd}_{k}$, $\boldsymbol{\eta}^{ad}_{k}$. During the initialization phase, both gyroscope and accelerometer biases are assumed to be approximately constant, i.e., $\mathbf{b}^{g}_{k}\approx\mathbf{b}^{g}$ and $\mathbf{b}^{a}_{k}\approx\mathbf{b}^{a}$.
IV Visual–Inertial Odometry
This section details the components of the proposed cost function. We first introduce the system state definition. We then describe the inertial and gravity-related terms that anchor the estimation to the IMU. Next, the visual reprojection factors derived from image measurements are presented. Finally, the non-rigid regularization terms, including elastic, viscous, and photometric components defined on the deformation graph, are introduced.
IV-A State definition
The system state is defined over a local temporal window spanning two consecutive keyframes, $\{t{-}1,t\}$. It comprises the rigid-body variables at both instants, namely orientation $\mathtt{R}_{t}\in\mathrm{SO}(3)$, velocity $\mathbf{v}_{t}\in\mathbb{R}^{3}$, and position $\mathbf{p}_{t}\in\mathbb{R}^{3}$, together with global inertial parameters, including the gyroscope and accelerometer biases $\mathbf{b}^{g},\mathbf{b}^{a}$ and the gravity direction $\hat{\mathbf{g}}\in\mathbb{S}^{2}$. We use the gravity direction representation introduced in [27], Sec. III–C. To account for non-rigid scene dynamics, the state is further augmented with a non-rigid substate $\boldsymbol{\xi}_{\mathrm{NR}}$, which collects the positions of all deformation nodes active within the window, stacked at times t1t{-}1 and $t$. Formally, the system state is given by
$\boldsymbol{\xi}=\big[\mathtt{R}_{t-1},\mathbf{v}_{t-1},\mathbf{p}_{t-1},\;\mathtt{R}_{t},\mathbf{v}_{t},\mathbf{p}_{t},\;\mathbf{b}^{g},\mathbf{b}^{a},\hat{\mathbf{g}},\;\boldsymbol{\xi}_{\mathrm{NR}}\big].$ (4)
In the following subsections, the corresponding residuals are developed.
IV-B IMU preintegration and gravity residuals
Inertial information is incorporated through IMU preintegration factors following the on-manifold formulation of [26]. Between two time instants $t_{i}$ and $t_{j}$, the relative motion is constrained by rotation, velocity, and position residuals,
$\mathbf{r}_{\Delta\mathtt{R}}$ $=\mathrm{Log}\!\left(\Delta\tilde{\mathtt{R}}_{t-1,t}^{\top}\mathtt{R}_{t-1}^{\top}\mathtt{R}_{t}\right),$ (5)
@@ IV-B IMU preintegration and gravity residuals
$\mathbf{r}_{\Delta\mathbf{v}}$ $=\mathtt{R}_{i}^{\top}\!\left(\mathbf{v}_{t}-\mathbf{v}_{t-1}-\mathbf{g}\Delta t_{t-1,t}\right)-\Delta\tilde{\mathbf{v}}_{t-1,t},$ (6)
$\mathbf{r}_{\Delta\mathbf{p}}=\mathtt{R}_{i}^{\top}\!\left(\mathbf{p}_{t}-\mathbf{p}_{t-1}-\mathbf{v}_{t-1}\Delta t_{t-1,t}-\tfrac{1}{2}\mathbf{g}\Delta t_{t-1,t}^{2}\right)-\Delta\tilde{\mathbf{p}}_{t-1,t}$ (7)
@@
Each residual is weighted by its corresponding covariance matrix, obtained during the preintegration process as described in [27], Sec. III–A. Additionally, a gravity residual is introduced to enforce consistency between the estimated accelerations and the gravity direction,
$\mathbf{r}_{\mathbf{g}}=\left(\frac{\mathbf{v}_{t}-\mathbf{v}_{t-1}}{\Delta t_{t-1,t}}-\frac{\mathtt{R}_{t-1}\Delta\mathbf{v}_{t-1,t}}{\Delta t_{t-1,t}}\right)-\|\mathbf{g}\|\,\hat{\mathbf{g}}$, (8)
@@
$\mathcal{L}_{\mathrm{imu}}^{t}=\|\mathbf{r}_{\Delta\mathtt{R}}\|_{{\boldsymbol{\Sigma}}_{\Delta\boldsymbol{\phi}}}+\|\mathbf{r}_{\Delta\mathbf{v}}\|_{{\boldsymbol{\Sigma}}_{\Delta\mathbf{v}}}+\|\mathbf{r}_{\Delta\mathbf{p}}\|_{{\boldsymbol{\Sigma}}_{\Delta\mathbf{p}}}+\|\mathbf{r}_{\mathbf{g}}\|_{{\boldsymbol{\Sigma}}_{\mathbf{g}}}$ (9)
IV-C Visual reprojection term
The visual contribution is modeled through a standard geometric reprojection error, which enforces multi-view consistency between observed image features and their predicted projections under the estimated camera motion. Specifically, the visual cost at time $t$ is defined as

### v2

@@ Abstract
Deformable scenes violate the rigidity assumptions underpinning classical visual–inertial odometry (VIO), often leading to over–fitting to local non–rigid motion or to severe camera pose drift when deformation dominates visual parallax. In this paper, we introduce DefVINS, the first visual–inertial odometry pipeline designed to operate in deformable environments. Our approach models the odometry state by decomposing it into a rigid, IMU-anchored component and a non-rigid scene warp represented by an embedded deformation graph. As a second contribution, we present VIMandala, the first benchmark containing real images and ground-truth camera poses for visual–inertial odometry in deformable scenes. In addition, we augment the synthetic Drunkard’s benchmark with simulated inertial measurements to further evaluate our pipeline under controlled conditions. We also provide an observability analysis of the visual–inertial deformable odometry problem, characterizing how inertial measurements constrain camera motion and render otherwise unobservable modes identifiable in the presence of deformation. This analysis motivates the use of IMU anchoring and leads to a conditioning-based activation strategy that avoids ill-posed updates under poor excitation. Experimental results on both in the synthetic Drunkard’s and our real VIMandala benchmarks show that DefVINS outperforms rigid visual–inertial and non-rigid visual odometry baselines. Our source code and data will be released upon acceptance.
@@ I Introduction
Visual–Inertial Odometry (VIO) and SLAM are foundational technologies driving modern mobile robotics and Augmented Reality (AR). The combination of a camera with an Inertial Measurement Unit (IMU) stands out as a particularly robust sensing modality. The IMU helps handle challenging conditions such as low-texture environments and rapid motions, while making roll and pitch directly observable [1]. By providing high-frequency measurements of linear acceleration and angular velocity, the IMU effectively constrains the short-term pose dynamics and improves the observability of orientation and scale components that remain ambiguous in pure monocular vision. In turn, the camera leverages exteroceptive landmarks from the environment that constrain its six-degrees-of-freedom motion for larger temporal windows, mitigating the inevitable proprioceptive drift [2, 3]. State-of-the-art VIO systems, such as VINS-Mono [3] and OKVIS [4], have demonstrated strong performance in static, rigid environments, establishing them as a de facto standard for ego-motion estimation.
@@
However, the efficacy of classical VIO relies fundamentally on the assumption of scene rigidity [5]. This assumption is violated in real-world scenarios, most notably when imaging deformable objects such as human bodies or clothing. When the underlying geometric model fails to account for non–rigid motion,the performance of VIO systems deteriorates significantly. In such cases, the estimation may suffer from early over–fitting of the rigid model to local non–rigid motion, or, more severely, from substantial drift when deformation-induced parallax dominates the inter-frame motion [6]. Consequently, achieving robust and accurate localization in highly dynamic and deformable environments remains an open and critical research challenge [7].
@@
Beyond modeling challenges, the transition from rigid to deformable environments fundamentally alters the observability properties of the estimation problem. Classical analyses of VIO observability have established that inertial sensing renders scale, gravity direction, and roll–pitch observable in rigid scenes. However, the implications of these properties in the presence of non-rigid motion remain largely unexplored. Deformation introduces additional latent degrees of freedom that may strongly couple with camera motion, leading to severe ill-conditioning and ambiguity when relying on visual information alone. In this context, inertial measurements play a crucial role by anchoring rigid-body dynamics over short time horizons, thereby constraining the solution space and reducing spurious correlations between ego-motion and non-rigid deformation. Motivated by this observation, our work presents an explicit observability analysis of the visual–inertial deformable odometry problem, showing that inertial measurements substantially improve conditioning and introduce structural constraints that are otherwise absent in purely visual formulations.
In this paper, we introduce DefVINS, the first visual–inertial odometry explicitly designed for deformable scenes. DefVINS incorporates an embedded deformation graph to model the non-rigid warping of the scene, while carefully decoupling the rigid, IMU-anchored state from the scene’s non-rigid degrees of freedom. Given the absence of benchmarks specifically targeting non-rigid VIO, we also contribute two new evaluation resources: (1) an extension of the synthetic Drunkard’s dataset augmented with inertial measurements, and (2) VIMandala, a novel real-world dataset featuring visual–inertial data and camera pose ground truth while observing a deformable mandala cloth. Experimental results on both datasets demonstrate that DefVINS consistently outperforms rigid visual–inertial methods as well as existing deformable visual odometry baselines.
@@ II Related Work
Visual–Inertial Odometry (VIO) is a well–established paradigm for robust ego–motion estimation in rigid environments through the tight fusion of visual and inertial measurements [8, 9, 10]. Inertial sensing provides high–rate motion constraints that resolve metric scale and stabilize roll and pitch, while visual observations offer global referencing and mitigate drift. The observability and consistency properties of such systems have been studied extensively under rigid–scene assumptions, showing that scale, gravity, velocity, and IMU biases become observable only under sufficient excitation and appropriate modeling choices [11, 12, 13]. These analyses, however, fundamentally assume that all observed features belong to a single rigid body.
When this assumption is violated, visual residuals become inconsistent and may bias the estimation or lead to divergence. Several approaches address dynamic scenes, composed of several rigid objects with independent motions, by detecting and suppressing dynamic elements via semantic segmentation or motion consistency checks [14, 15, 16]. While effective under moderate dynamic content, these methods rely on rigid scene parts for camera motion estimation, which limit their applicability in scenes dominated by non–rigid motion. Other approaches are able to track dynamic content [17, 18, 19, 20] and may be more resilient to the lack of rigid scene parts. However, they also model the scene as a set of rigidly moving objects, limiting certain applications.
A complementary line of work explicitly models fully non–rigid scene geometry, targeting among others human or medical reconstruction. DynamicFusion [21], DefSLAM [22], and NR–SLAM [23] represent deformation using embedded deformation graphs or learned warp fields, enabling compelling reconstructions. However, DynamicFusion relies on RGB-D sensors, which are rarely found in some setups. And DefSLAM and NR–SLAM rely exclusively on visual cues, lack inertial anchoring, and do not preserve metric scale, resulting in a strong coupling between camera motion and scene deformation and, consequently, ill–conditioned pose estimation under strong non–rigid motion.
@@
To the best of our knowledge, no existing method jointly (i) integrates an explicit deformation model within a visual–inertial odometry pipeline, and (ii) maintains a rigid, IMU–anchored reference to ensure metric consistency. DefVINS addresses these limitations by embedding a deformation graph within a visual–inertial optimization framework and by explicitly analyzing the observability properties of the resulting system, demonstrating that inertial constraints significantly improve conditioning even under strong non–rigid motion.
III Visual–Inertial Odometry
III-A State definition
[…]
$\small\boldsymbol{\xi}=\mathrm{stack}\Big(\{\mathtt{R}_{\tau},\mathbf{v}_{\tau},\mathbf{t}_{\tau}\}_{\tau\in\mathcal{W}},\;\mathbf{b}^{g},\mathbf{b}^{a},\hat{\mathbf{g}},\;\{\mathbf{x}_{i}^{\tau}\}_{i\in\mathcal{D},\tau\in\mathcal{W}}\Big)$ (1)
III-B IMU preintegration and gravity residuals
The IMU kinematics between the time instants of two consecutive keyframes, $t_{\tau}$ and $t_{\tau+1}$, are as follows
@@ III-B IMU preintegration and gravity residuals
$\mathtt{R}_{\tau+1}$ $=\mathtt{R}_{\tau}\prod_{t_{k}\in\left[t_{\tau},t_{\tau+1}\right)}\mathrm{Exp}\!\left(\left(\tilde{\boldsymbol{\omega}}_{k}-\mathbf{b}^{g}-\boldsymbol{\eta}^{gd}_{k}\right)\,\Delta t\right),$ (2)
$\mathbf{v}_{\tau+1}$ $=\mathbf{v}_{\tau}+\mathbf{g}\,\Delta T+\sum_{t_{k}\in\left[t_{\tau},t_{\tau+1}\right)}\mathtt{R}_{k}\big(\tilde{\mathbf{a}}_{k}-\mathbf{b}^{a}-\boldsymbol{\eta}^{ad}_{k}\big)\,\Delta t,$ (3)
$\mathbf{t}_{\tau+1}$ $=\mathbf{t}_{\tau}+\sum_{t_{k}\in\left[t_{\tau},t_{\tau+1}\right)}\left(\mathbf{v}_{k}\Delta t+\tfrac{1}{2}\mathbf{g}\,\Delta t^{2}\right)$
@@
$\quad+\sum_{t_{k}\in\left[t_{\tau},t_{\tau+1}\right)}\frac{1}{2}\mathtt{R}_{k}\left(\tilde{\mathbf{a}}_{k}-\mathbf{b}^{a}-\boldsymbol{\eta}^{ad}_{k}\right)\,\Delta t^{2},$ (4)
where $\Delta t$ is the IMU sampling period and $\tilde{\boldsymbol{\omega}}_{k}$ and $\tilde{\mathbf{a}}_{k}$ correspond to the gyroscope and accelerometer readings at $t_{k}\in\left[t_{\tau},t_{\tau+1}\right]$, which are affected by biases $\mathbf{b}^{g}$, $\mathbf{b}^{a}$ and additive noise $\boldsymbol{\eta}^{gd}_{k}$, $\boldsymbol{\eta}^{ad}_{k}$. Between time instants $t_{\tau}$ and $t_{\tau+1}$, the relative motion is constrained by rotation, velocity, and position residuals,
$\mathbf{r}_{\Delta\mathtt{R}}$ $=\mathrm{Log}\left(\Delta\tilde{\mathtt{R}}_{\tau,\tau+1}^{\top}\mathtt{R}_{\tau}^{\top}\mathtt{R}_{\tau+1}\right)$ (5)
$\mathbf{r}_{\Delta\mathbf{v}}$ $=\mathtt{R}_{\tau}^{\top}(\mathbf{v}_{\tau+1}-\mathbf{v}_{\tau}-\mathbf{g}\Delta T)-\Delta\tilde{\mathbf{v}}_{\tau,\tau+1}$ (6)
$\mathbf{r}_{\Delta\mathbf{t}}$ $=\mathtt{R}_{\tau}^{\top}(\mathbf{t}_{\tau+1}-\mathbf{t}_{\tau}-\mathbf{v}_{\tau}\Delta T-\tfrac{1}{2}\mathbf{g}\Delta T^{2})-\Delta\tilde{\mathbf{t}}_{\tau,\tau+1}$ (7)
@@
where $\Delta\tilde{\mathtt{R}}_{\tau,\tau+1}$, $\Delta\tilde{\mathbf{v}}_{\tau,\tau+1}$ and $\Delta\tilde{\mathbf{t}}_{\tau,\tau+1}$ are preintegrated values from the IMU readings [26]. Each of them is weighted by its corresponding covariance matrix ${\boldsymbol{\Sigma}}_{\Delta\boldsymbol{\phi}}$}, ${\boldsymbol{\Sigma}}_{\Delta\mathbf{v}}$ and ${\boldsymbol{\Sigma}}_{\Delta\mathbf{t}}$ obtained during the preintegration process as described in [27], Sec. III–A. Additionally, a gravity residual is introduced to enforce consistency between the estimated accelerations and the gravity direction,
$\mathbf{r}_{\mathbf{g}}=\left(\frac{\mathbf{v}_{\tau+1}-\mathbf{v}_{\tau}}{\Delta T}-\frac{\mathtt{R}_{\tau}\Delta\mathbf{v}_{\tau,\tau+1}}{\Delta T}\right)-\|\mathbf{g}\|\,\hat{\mathbf{g}}$, (8)
@@
$\mathcal{L}_{\mathrm{imu}}^{\tau}=\|\mathbf{r}_{\Delta\mathtt{R}}\|_{{\boldsymbol{\Sigma}}_{\Delta\boldsymbol{\phi}}}+\|\mathbf{r}_{\Delta\mathbf{v}}\|_{{\boldsymbol{\Sigma}}_{\Delta\mathbf{v}}}+\|\mathbf{r}_{\Delta\mathbf{t}}\|_{{\boldsymbol{\Sigma}}_{\Delta\mathbf{t}}}+\|\mathbf{r}_{\mathbf{g}}\|_{{\boldsymbol{\Sigma}}_{\mathbf{g}}}$ (9)
III-C Reprojection error
Each tracked feature will contribute with its reprojection error, that will be aggregated for the keyframe at time $t_{\tau}$ as